Problem 2. Sender Utilization (20 pt.)

  1. Assume stop-and-wait, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization.
  1. Assume pipelining and 20 packets can be sent without ACK, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization.
Answer #1

a)

Given that Bandwidth = 100 Mbps, Propagation delay(Tp) = 10 ms and packet length = 10000 bits

Transmission time(Tt) = packet length/Bandwidth = ( 10, 000)/(100 * 10^6) = 0.0001 seconds = 0.1 ms

So, sender utilization = Tt/(Tt + 2Tp) = 0.1/(0.1 + 2*10) = 0.1/20.1 = 0.49 % utilization

b)

Given that 20 packets (N)  can be sent without ACK using pipeling. Therefore, N*Tt time for transmitting N packets.

So, the sender utilization is (N*Tt)/(Tt + 2*Tp) = (20*0.1)/20.1 = 9.95 % utilization.